OpenStudy (anonymous):

solve the system of linear equations: 3x-2y = 3 4x + 3y = 38

OpenStudy (anonymous):

7x+3y=35

OpenStudy (anonymous):

find the value of xy.... sorry i forgot to right that down

OpenStudy (anonymous):

let x=0 7(0)+3y=35 3y=35 divide both side by 3 so y = 35/3 7x+3(35/3)=35 7x+35=35 7x=35-35 7x=0 divide both side by 7 x=0 x=0 and y=35/3